# Debating Truth: Debate-driven Claim Verification with Multiple Large Language Model Agents

### v2

@@ 5.1. Experimental Setup
To evaluate DebateCV, we use AVeriTeC (averitec), the state-of-the-art (SOTA) real-world claim verification benchmark. AVeriTeC consists of English textual claims collected from 50 fact-checking agencies and annotated by human fact-checkers, which addresses key limitations of previous datasets, such as evidence leakage and insufficiency (defame; fact2fiction). Each claim in AVeriTeC is paired with a knowledge base containing both relevant and potentially distracting evidence collected via Google Search, which simulates open-web retrieval while ensuring reproducibility (averitec_challenge). To systematically evaluate our method under different evidence quality conditions, we consider three distinct evidence settings:
@@
  • Retrieved (H): Evidence retrieved using the HerO framework (hero), which achieved the best evidence retrieval performance in the AVeriTeC challenge (averitec_challenge). This condition simulates realistic end-to-end fact-checking pipelines where retrieval may introduce irrelevant or misleading evidence. We use the pre-retrieved evidence provided by hero. [^8] https://github.com/ssu-humane/HerO/
  • Retrieved (I): Evidence retrieved using the InFact framework (infact), which ranked second in the challenge above. This additional condition allows us to assess the robustness of DebateCV across different evidence retrieval strategies. We use the official InFact implementation with GPT-4o-Mini-2024-07-18 as the backbone. [^9] https://github.com/multimodal-ai-lab/DEFAME/tree/infact
@@
Single-agent baselines prompt a single LLM to use the provided evidence to predict the verdict of the claims. SOTA methods employ the chain-of-thought (CoT) (cot) prompting strategy to enhance the reasoning capabilities of zero-shot LLMs. Specifically, we evaluate two such prompting strategies: CoT (yang_and_rocha), and InFact (infact). In addition to zero-shot methods, we also include (3) HerO (hero), which conducts supervised fine-tuning (SFT) on the training split of the AVeriTeC dataset to further improve the performance of single LLMs.
Multi-agent baselines include: (1) Majority: fact_audit propose employing three independent agents to verify each claim, with the final verdict determined by majority voting among the agents. In our implementation, each agent uses the CoT baseline, and we adopt the majority voting prompt from the official implementation of fact_audit. [^11] https://github.com/DanielLin97/FACT-AUDIT (2) Cooperation: This baseline utilises the CAMEL framework (camel), which simulates collaborative task-solving through structured interactions between an AI user and an AI assistant. We leverage the official implementation of CAMEL and use the prompt of the CoT baseline as the task description. [^12] https://github.com/camel-ai/camel
@@
We evaluate the baselines using four LLMs: two proprietary models, GPT-4o-2024-05-13 (GPT-4o) and GPT-4o-Mini-2024-07-18 (GPT-4o-Mini), and two open-source models, Llama-3.1-8B (llama) and Qwen-2.5-7B (qwen). For HerO, which requires supervised fine-tuning, we restrict our experiments to the open-source models due to the unavailability of trainable parameters for GPT-4o and GPT-4o-Mini. In DebateCV, the Debaters employ GPT-4o-Mini to balance cost efficiency with strong performance. [^13] Sec. 5.5 validates that DebateCV is robust to alternative Debaters. The Moderator, which produces the final verdict and justification, uses Llama-3.1-8B and Qwen-2.5-7B to enable a direct comparison with HerO (hero). This configuration ensures a comprehensive evaluation across diverse LLM architectures.
We set the maximum number of generated tokens to 512, temperature to 0.7, and top_p to 1.0. For post-training, we employ LoRA (lora) with rank 128 and alpha 256 and the Adam optimizer with a learning rate of 2e-5 for 2 epochs. These hyper-parameters are identical to those used by hero to ensure a fair comparison.
@@
We employ two complementary metrics to evaluate the methods. (1) Accuracy (Acc.): The proportion of correctly classified claims across all test instances, which measures an overall performance regardless of evidence sufficiency. (2) AVeriTeC Score (Aver.): Accuracy conditioned on evidence sufficiency (averitec), which excludes correctly classified claims that lack sufficient supporting evidence, thereby avoiding score inflation due to lucky guesses. This metric determines evidence sufficiency using the METEOR score (meteor) between retrieved evidence $\hat{\mathcal{E}}$ and golden evidence $\mathcal{E}$, denoted as f(^,)f($\hat{\mathcal{E}}$$\mathcal{E}$). Claims are deemed to have sufficient evidence if $f(\hat{\mathcal{E}},\mathcal{E})\geq 0.25$ (averitec). [^14] Note that under the golden evidence condition, the AVeriTeC score equals accuracy; therefore, we report only accuracy for this setting. For both metrics, we follow complex_method to use paired bootstrap tests to assess statistical significance based on at least five repeated trials. Results with $p\leq 0.05$ compared to the strongest baseline (HerO) are marked with $\dagger$”.
@@ 5.2. Comparison with Existing Methods
Table 1. Claim verification performance across varying evidence conditions. Results are reported as percentages (%). The best result in each column is bolded; the second-best is underlined.
@@
Zero-shot single-agent baselines (CoT and InFact) and multi-agent approaches (Majority and Cooperation) both exhibit limited robustness under realistic retrieved evidence settings. For example, when using GPT-4o as the backbone model, CoT’s accuracy drops from 81.6% with golden evidence to 63.8% in the Retrieved (H) setting by 17.8%. Multi-agent methods leverage collective intelligence but yield only modest improvements over single-agent counterparts; for instance, Majority achieves merely a 0.6 point gain over CoT with GPT-4o under both Retrieved (H) and Retrieved (I) conditions, while Cooperation, which simulates user-assistant dialogue for evidence analysis, only improves with strong LLMs like GPT-4o but can perform worse with smaller models, such as Qwen-2.5-7B in Retrieved (H), where accuracy drops from 62.6% to 60.8%. The results underscore the critical challenge of handling noisy retrieved evidence in end-to-end fact-checking pipelines. The limitations of these methods stem from the absence of cross-validation mechanisms in existing methods. Single-agent reasoning follows a linear chain of thought, rendering it vulnerable to misleading or incomplete evidence (for example, those introduced in Sec. 1 and 5.2). Multi-agent baselines also lack rigorous adversarial critique, reinforcing erroneous consensus rather than correcting it.
Our debate-driven claim verification framework, DebateCV, consistently outperforms all baselines across every evidence condition. For instance, when powered by Llama-3.1-8B, DebateCV achieves the highest scores: 83.4% under the golden evidence setting, 72.8% / 54.4% under Retrieved (H), and 73.6% / 54.8% under Retrieved (I). These results demonstrate that our debate-driven approach significantly surpasses the prior state-of-the-art HerO method (hero), delivering absolute gains of at least 2.0% under ideal golden evidence and 2.6–5.8% under realistic retrieved evidence conditions. This superiority arises from the adversarial cross-agent critique inherent in the debate process, which rigorously evaluates evidence and effectively mitigates the influence of misleading or ambiguous information (see the qualitative case study in Sec. 5.2 for empirical support for this finding). The DebateCV variant using Qwen-2.5-7B as its backbone also delivers strong performance, securing the second-best results across all settings: 82.0% (Golden), 72.6% / 54.6% (Retrieved (H)), and 72.4% / 54.6% (Retrieved (I)), respectively. This consistency underscores the robustness of our framework across different LLM backbones.
@@
In addition to accuracy metrics, we conduct a Mean Opinion Score (MOS) evaluation of justification quality based on the criteria proposed by warren2025show, who identified three key dimensions of justification quality through interviews with fact-checking experts across five continents: (1) Evidence Used: clarity and comprehensiveness in presenting supporting evidence; (2) Sources of Uncertainty: appropriate acknowledgment of evidence limitations; and (3) Reasoning Pathway: explicit, logical connections from evidence to conclusion.
@@
Table 3. The claim verification results of DebateCV. Responses are condensed for brevity while preserving original semantics. Erroneous judgements are in red, accurate judgements are in green, and adversarial arguments in DebateCV are in blue.
@@
(1) Claim/evidence misinterpretation. Regarding Claim 31 on Amy Barrett’s confirmation, CoT incorrectly conflates her confirmation date (October 26) with her assumption of office (October 27), yielding an incorrect “Conflicting Evidence/Cherry-picking” verdict. HerO correctly identifies the confirmation date but offers only a superficial justification, failing to address the potential ambiguity between confirmation and assumption dates. In contrast, DebateCV’s debate structure enables nuanced interpretation: the Affirmative Debater supports the claim using confirmation evidence, while the Negative Debater challenges it by interpreting “confirmed” as referring to the assumption of duties. This exchange allows the Moderator to disambiguate the claim’s true intent regarding the confirmation act itself and to issue a correct Supported verdict.
@@
Figure 2. Claim verification performance across debated rounds.
Table 4. False positive rates for neutral verdicts (NEE and CEC).
@@ 5.3. Ablation Study
The results in Table 1 reveal complementary roles for each component. Removing $\mathcal{C}_{\text{error}}$ causes substantial performance drops across all conditions. For Llama-3.1-8B, accuracy decreases from 83.4% to 76.4% with Golden evidence, from 72.8% to 63.0% under Retrieved (H), and from 73.6% to 68.2% with Retrieved (I). Qwen-2.5-7B exhibits similar degradation patterns. Removing $\mathcal{C}_{\text{correct}}$, however, shows divergent effects depending on the evidence condition: it preserves competitive performance when using retrieved evidence but incurs more substantial performance degradation under golden evidence. For instance, Llama-3.1-8B attains 70.8% accuracy under Retrieved (H), which is only 2.4% below that of the full model. Similarly, its AVeriTeC score also remains almost unaffected (dropping only 0.4-1.6% in all setups). In contrast, its accuracy drops more markedly under golden evidence, falling to 76.8% from the full model’s 83.4%. The results show that the correct samples $\mathcal{C}_{\text{correct}}$ are only important for high-quality evidence scenarios, while corrected samples $\mathcal{C}_{\text{error}}$ are crucial for robust performance under noisy evidence conditions, highlighting the importance of our error correction mechanism. Incorporating both data components achieves optimal performance across varying evidence quality.
@@
As shown in Table 1, our Debate-SFT method yields substantial improvements in the Moderator’s performance compared to its zero-shot counterpart. For instance, under Retrieved (H), the w/o Debate-SFT variant exhibits significant performance degradation, with accuracy dropping from 72.8% to 53.8% when using Llama-3.1-8B and from 72.6% to 54.2% when using Qwen-2.5-7B. The results highlight the risks of using zero-shot LLM agents, such as those in (han2025beyond), for this task.
To further interpret these improvements, Fig. 2 compares DebateCV performance with and without Debate-SFT, using Llama-3.1-8B as Moderators across debated rounds under the Retrieved (H) condition. It is observed that the zero-shot Moderator resolves most claims in the first debate round, but its error rate rises in subsequent rounds. This is likely due to the inherent complexity of claims requiring extended debate, compounded by known LLM performance degradation in multi-turn dialogues in various tasks (llm_lost; tcss_stance). In contrast, the Moderator post-trained by Debate-SFT sustains higher accuracy across all rounds, which shows enhanced decision-making for both simple claims resolvable in a single round and more challenging ones.
Further, we observe that the zero-shot Moderator frequently defaults to neutral verdicts, i.e., Not Enough Evidence (NEE) and Conflicting Evidence/Cherry-picking (CEC), even when sufficient evidence exists to definitively support or refute a claim. This may stem from conformity bias in LLMs (conform1; conform2), where the Moderator avoids contradicting either Debater. For example, in Table 4, the zero-shot Moderator yields high false positive rates of 25.0% for NEE and 6.3% for CEC. Nonetheless, Debate-SFT dramatically reduces the false positive rates to 0.4% for NEE and 1.1% for CEC, respectively, highlighting its benefits in mitigating unwarranted neutrality and generating decisive verdicts.
@@ 5.4. Human Evaluation on Synthetic Data
To validate the synthetic data in the SynDeC dataset, i.e., the debate recordings and the corrected justifications for $\mathcal{C}_{\text{error}}$, we recruited the same participants in the human evaluation in Sec. 5.2 to evaluate the quality of the synthetic debate recordings and justifications for 50 randomly selected samples. Following the evaluation protocol of nan2024let, we asked the participants to rate each sample on a three-point Likert scale (“not likely,” “likely,” or “very likely”) in response to the following questions:
@@
    Question 2 (for Justifications): “To what extent does this justification accurately summarize and reflect the debated arguments?
@@
As shown in Fig. 3, both the generated debates and justifications received high ratings: 96.2% of debates were rated as “likely” or “very likely” for Question 1, and over 94.8% of justifications received the same ratings for Question 2. Inter-annotator agreement was high, with Krippendorff’s $\alpha=0.83$ for debate quality (excellent) and $\alpha=0.71$ for justification quality (substantial), respectively. These results indicate that our Debaters consistently generate substantive discussions to support their reasoning, and that the resulting justifications effectively distill the arguments presented in the debates, all without hallucinations.

### v3

@@ 5.1. Experimental Setup
To evaluate DebateCV, we use AVeriTeC (Schlichtkrull et al., 2024b), the state-of-the-art (SOTA) real-world claim verification benchmark, which addresses the evidence leakage and insufficiency issues in prior datasets (Braun et al., 2025; He et al., 2026) like FEVER (Thorne et al., 2018). AVeriTeC consists of English textual claims collected from 50 fact-checking agencies and annotated by human fact-checkers. Each claim in AVeriTeC is paired with a knowledge base containing both relevant and potentially distracting evidence collected via Google Search, which simulates open-web retrieval while ensuring reproducibility (Schlichtkrull et al., 2024a). To evaluate our method under different evidence quality conditions, we consider three distinct evidence settings:
@@
  • Retrieved (H): Evidence retrieved using the HerO framework (Yoon et al., 2024), which achieved the best evidence retrieval performance in the AVeriTeC challenge (Schlichtkrull et al., 2024a). This condition simulates realistic end-to-end fact-checking pipelines where retrieval may introduce irrelevant or misleading evidence. We use the pre-retrieved evidence provided by Yoon et al. (2024). [^8] https://github.com/ssu-humane/HerO/
  • Retrieved (I): Evidence retrieved using the InFact framework (Rothermel et al., 2024), which ranked second in the challenge above. This additional condition allows us to assess the robustness of DebateCV across different evidence retrieval strategies. We use the official InFact implementation with GPT-4o-Mini-2024-07-18 as the backbone. [^9] https://github.com/multimodal-ai-lab/DEFAME/tree/infact
@@
Single-agent baselines prompt a single LLM to use the provided evidence to predict the verdict of the claims. SOTA methods employ the chain-of-thought (CoT) prompting strategy (Wei et al., 2022) to enhance the reasoning capabilities of zero-shot LLMs. Specifically, we evaluate two such prompting strategies: CoT (Yang and Rocha, 2024), and InFact (Rothermel et al., 2024). In addition to zero-shot methods, we also include (3) HerO (Yoon et al., 2024), which conducts supervised fine-tuning (SFT) on the training split of the AVeriTeC dataset to further improve the performance of single LLMs.
Multi-agent baselines include: (1) Majority: Lin et al. (2025) propose employing three independent agents to verify each claim, with the final verdict determined by majority voting among the agents. In our implementation, each agent uses the CoT baseline, and we adopt the majority voting prompt from the official implementation of Lin et al. (2025). [^11] https://github.com/DanielLin97/FACT-AUDIT (2) Cooperation: This baseline utilises the CAMEL framework (Li et al., 2023a), which simulates collaborative task-solving through structured interactions between an AI user and an AI assistant. We leverage the official implementation of CAMEL and use the prompt of the CoT baseline as the task description. [^12] https://github.com/camel-ai/camel
@@
We evaluate the baselines using four LLMs: two proprietary models, GPT-4o-2024-05-13 (GPT-4o) and GPT-4o-Mini-2024-07-18 (GPT-4o-Mini), and two open-source models, Llama-3.1-8B (Meta, 2024) and Qwen-2.5-7B (Qwen, 2025). For HerO, which requires supervised fine-tuning, we restrict our experiments to the open-source models due to the unavailability of trainable parameters for GPT-4o and GPT-4o-Mini. In DebateCV, the Debaters employ GPT-4o-Mini to balance cost efficiency with strong performance. [^13] Sec. 5.5 validates that DebateCV is robust to alternative Debaters. The Moderator, which produces the final verdict and justification, uses Llama-3.1-8B and Qwen-2.5-7B to enable a direct comparison with HerO (Yoon et al., 2024). This configuration ensures a comprehensive evaluation across diverse LLM architectures.
We set the maximum number of generated tokens to 512, temperature to 0.7, and top_p to 1.0. For post-training, we employ LoRA with rank 128 and alpha 256 and the Adam optimizer with a learning rate of 2e-5 for 2 epochs. These hyper-parameters are identical to those used by Yoon et al. (2024) to ensure a fair comparison.
@@
We employ two complementary metrics to evaluate the methods. (1) Accuracy (Acc.): The proportion of correctly classified claims across all test instances, which measures an overall performance regardless of evidence sufficiency. (2) AVeriTeC Score (Aver.): Accuracy conditioned on evidence sufficiency (Schlichtkrull et al., 2024b), which excludes correctly classified claims that lack sufficient supporting evidence, thereby avoiding score inflation due to lucky guesses. We follow Schlichtkrull et al. (2024b) to use the METEOR score (Banerjee and Lavie, 2005) to assess evidence sufficiency between retrieved evidence $\hat{\mathcal{E}}$ and golden evidence $\mathcal{E}$, denoted as f(^,)f($\hat{\mathcal{E}}$$\mathcal{E}$). Claims are deemed to have sufficient evidence if $f(\hat{\mathcal{E}},\mathcal{E})\geq 0.25$ (Schlichtkrull et al., 2024b). [^14] Note that under the golden evidence condition, the AVeriTeC score equals accuracy; therefore, we report only accuracy for this setting. For both metrics, we follow Chen et al. (2024) to use paired bootstrap tests to assess statistical significance based on at least five repeated trials. Results with $p\leq 0.05$ compared to the strongest baseline (HerO) are marked with $\dagger$”.
@@ 5.2. Comparison with Existing Methods
Table 1. Claim verification performance across evidence conditions (reported as percentages). The best result in each column is bolded; the second-best is underlined.
@@
Zero-shot single-agent baselines (CoT and InFact) and multi-agent approaches (Majority and Cooperation) both exhibit limited robustness under realistic retrieved evidence settings. For example, when using GPT-4o as the backbone model, CoT’s accuracy drops from 81.6% with golden evidence to 63.8% in the Retrieved (H) setting by 17.8%. Multi-agent methods leverage collective intelligence but yield only modest improvements over single-agent counterparts; for instance, Majority achieves merely a 0.6 point gain over CoT with GPT-4o under both Retrieved (H) and Retrieved (I) conditions, while Cooperation, which simulates user-assistant dialogue for evidence analysis, only improves with strong LLMs like GPT-4o but can perform worse with smaller models, such as Qwen-2.5-7B in Retrieved (H), where accuracy drops from 62.6% to 60.8%. The results underscore the critical challenge of handling noisy retrieved evidence in end-to-end fact-checking pipelines. These limitations stem from the absence of cross-validation mechanisms in existing approaches. Single-agent reasoning follows a linear chain of thought, rendering it vulnerable to misleading or incomplete evidence (for example, those introduced in Sec. 1 and 5.2). Multi-agent baselines also lack rigorous adversarial critique, reinforcing erroneous consensus rather than correcting it.
Our debate-driven claim verification framework, DebateCV, consistently outperforms all baselines across every evidence condition. For instance, when powered by Llama-3.1-8B, DebateCV achieves the highest scores: 83.4% under the golden evidence setting, 72.8% / 54.4% under Retrieved (H), and 73.6% / 54.8% under Retrieved (I). These results demonstrate that our debate-driven approach significantly surpasses the prior state-of-the-art HerO method (Yoon et al., 2024), delivering absolute gains of at least 2.0% under ideal golden evidence and 2.6–5.8% under realistic retrieved evidence conditions. This superiority arises from the adversarial cross-agent critique inherent in the debate process, which rigorously evaluates evidence and effectively mitigates the influence of misleading or ambiguous information (see the qualitative case study in Sec. 5.2 for empirical support for this finding). The DebateCV variant using Qwen-2.5-7B as its backbone also delivers strong performance, securing the second-best results across all settings: 82.0% (Golden), 72.6% / 54.6% (Retrieved (H)), and 72.4% / 54.6% (Retrieved (I)), respectively. This underscores the robustness of our framework across different LLM backbones.
@@
Beyond accuracy, we conduct a Mean Opinion Score (MOS) evaluation (Streijl et al., 2016) of justification quality based on the criteria proposed by Warren et al. (2025), who identified three key dimensions of justification quality by interviews with fact-checking experts across five continents: (1) Evidence Used: clarity and comprehensiveness in presenting supporting evidence; (2) Sources of Uncertainty: appropriate acknowledgment of evidence limitations; and (3) Reasoning Pathway: explicit, logical connections from evidence to conclusion.
@@
Table 3. The claim verification results of DebateCV. Responses are condensed for brevity while preserving original semantics. Erroneous judgements are in purple, accurate judgements are in green, and adversarial arguments in DebateCV are in blue.
@@
(1) Claim/evidence misinterpretation. For Claim 31 on Amy Barrett’s confirmation, CoT incorrectly conflates her confirmation date (October 26) with her assumption of office (October 27), yielding an incorrect “Conflicting Evidence/Cherry-picking” verdict. HerO correctly identifies the confirmation date but offers only a superficial justification, failing to address the potential ambiguity between confirmation and assumption dates. In contrast, DebateCV’s debate structure enables nuanced interpretation: the Affirmative Debater supports the claim using confirmation evidence, while the Negative Debater challenges it by interpreting “confirmed” as referring to the assumption of duties. This exchange allows the Moderator to disambiguate the claim’s true intent regarding the confirmation act itself and to issue a correct Supported verdict.
@@
Figure 2. Claim verification performance across rounds.
Table 4. False positive rates for neutral verdicts.
@@ 5.3. Ablation Study
Table 1 reveals complementary roles for each component. Removing $\mathcal{C}_{\text{error}}$ causes substantial performance drops across all conditions. For Llama-3.1-8B, accuracy decreases from 83.4% to 76.4% with Golden evidence, from 72.8% to 63.0% under Retrieved (H), and from 73.6% to 68.2% with Retrieved (I). Qwen-2.5-7B exhibits similar degradation patterns. Removing $\mathcal{C}_{\text{correct}}$, however, shows divergent effects depending on the evidence condition: it preserves competitive performance when using retrieved evidence but incurs more substantial performance degradation under golden evidence. For instance, Llama-3.1-8B attains 70.8% accuracy under Retrieved (H), which is only 2.4% below that of the full model. Similarly, its AVeriTeC score also remains almost unaffected (dropping only 0.4-1.6% in all setups). In contrast, its accuracy drops more markedly under golden evidence, falling to 76.8% from the full model’s 83.4%. The results show that the correct samples $\mathcal{C}_{\text{correct}}$ are only important for high-quality evidence scenarios, while corrected samples $\mathcal{C}_{\text{error}}$ are crucial for robust performance under noisy evidence conditions, highlighting the importance of our error correction mechanism. Incorporating both data components achieves optimal performance across all conditions.
@@
As shown in Table 1, our Debate-SFT method yields substantial improvements in the Moderator’s performance compared to its zero-shot counterpart. For instance, under Retrieved (H), the w/o Debate-SFT variant exhibits significant performance degradation, with accuracy dropping from 72.8% to 53.8% when using Llama-3.1-8B and from 72.6% to 54.2% when using Qwen-2.5-7B. The results highlight the risks of using zero-shot LLM agents, such as those in (Han et al., 2025a), for this task.
To further interpret these improvements, Fig. 2 compares DebateCV performance with and without Debate-SFT, using Llama-3.1-8B as Moderators across debated rounds under the Retrieved (H) condition. It is observed that the zero-shot Moderator resolves most claims in the first debate round, but its error rate rises in subsequent rounds. This is likely due to the inherent complexity of claims requiring extended debate, compounded by known LLM performance degradation in multi-turn dialogues in various tasks (Laban et al., 2025; Li et al., 2023b, c). In contrast, the Moderator post-trained by Debate-SFT sustains higher accuracy across all rounds, which shows enhanced decision-making for both simple claims resolvable in a single round and more challenging ones.
Further, we observe that the zero-shot Moderator frequently defaults to neutral verdicts, i.e., Not Enough Evidence (NEE) and Conflicting Evidence/Cherry-picking (CEC), even when sufficient evidence exists to definitively support or refute a claim. This may stem from conformity bias in LLMs (Weng et al., 2025; Zhang et al., 2024), where the Moderator avoids contradicting either Debater. For example, in Table 4, the zero-shot Moderator yields high false positive rates of 25.0% for NEE and 6.3% for CEC. Nonetheless, Debate-SFT dramatically reduces the false positive rates to 0.4% for NEE and 1.1% for CEC, respectively, highlighting its benefits in mitigating unwarranted neutrality and generating decisive verdicts.
@@ 5.4. Human Evaluation on Synthetic Data
To validate the synthetic data in the SynDeC dataset, i.e., the debate recordings and the corrected justifications for $\mathcal{C}_{\text{error}}$, we recruited the same participants in the human evaluation in Sec. 5.2 to evaluate the quality of the synthetic debate recordings and justifications for 50 randomly selected samples. Following the evaluation protocol of Nan et al. (2024), we asked the participants to rate each sample on a three-point Likert scale (“not likely,” “likely,” or “very likely”) in response to the following questions:
@@
    Question 2 (for Justifications): “To what extent does this justification accurately summarize and reflect the debate?
@@
As shown in Fig. 3, both the generated debates and justifications received high ratings: 96.2% of debates were rated as “likely” or “very likely” for Question 1, and over 94.8% of justifications received the same ratings for Question 2. Inter-annotator agreement was high, with Krippendorff’s $\alpha=0.83$ for debate quality (excellent) and $\alpha=0.71$ for justification quality (substantial), respectively. These results indicate that our Debaters consistently generate substantive discussions to support their reasoning, and that the resulting justifications effectively distil the arguments presented in the debates, all without hallucinations.